OpenStudy (anonymous):

If given the orbital period and radius of Venus, how would one determine the mass of the Sun?

OpenStudy (anonymous):

Orbital Period as in revolution period

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

And mean orbital radius

OpenStudy (anonymous):

is orbital peroid and radius both given

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

use kepler's third law

OpenStudy (anonymous):

we have radial velocity=2(pi)r/T centrifugal force = gravitational force therefore,GMm/r^2=mv^2/r [m=mass of the planet,G=universal gravitational constant,M=mass of the sun,v=circular velocity] the equation turns into GM/r = v^2 so mass of the sun M=r.v^2/G now use v=2.(pi).r